Consider a linear regression model where y represents the response variable, x is a quantitative explanatory variable, and d is a dummy variable. The model is estimated as  yˆy^  = 14.4 + 4.6x − 3.1d.

a. Interpret the dummy variable coefficient.

  • Intercept shifts down by 3.1 units as d changes from 0 to 1.

  • Slope shifts down by 3.1 units as d changes from 0 to 1.

  • Intercept shifts up by 3.1 units as d changes from 0 to 1.

  • Slope shifts up by 3.1 units as d changes from 0 to 1.

b. Compute yˆy^ for x = 3 and d = 1. (Round your answer to 1 decimal place.)

c. Compute yˆy^  for x = 3 and d = 0. (Round your answer to 1 decimal place.)
Answer #1

Ans:

a)

Intercept shifts down by 3.1 units as d changes from 0 to 1.

b)

when x=3 and d=1

y' = 14.4 + 4.6*3 − 3.1*1=25.1

c)

when x=3 and d=0

y' = 14.4 + 4.6*3 − 3.1*0=28.2